The coil of an electric generator rotates at a fre
Answer #1

f = 60 Hz

V = 120 V

A = 0.0030 m2

N = 500

We now for a generator,

V = NBAW/sqrt(2) <------- rms value of emf

where, W = 2*pi*f

So, V = 500*B*0.0030*(2*pi*60)/sqrt(2) = 120

So, B = 0.3 T <--------answer